Shift report 20000720 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Jul 20 2000 - 00:26:44 EDT

  • Next message: http account: "Shift report 20000720 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Jul 20 00:26:43 2000, Shift: 16:00-24:00
    Supervisor: fv, with: HI
    Log book#3 pages 30 - 30
    
    Shift Summary:
    ------------------------------------------------------------
    RHIC worked to improve tune at injection;
    A couple of period 3*40 minutes at inection; 
    and a single ramp attempt. Otherwise booster extraction
    problem(s).
    
    No problems reported:
    - a curiousty . If the DAQ is put in pulser mode  as seperate from
    pedestal or data. The triggerbits are ignored all 1.
    
    Achievements:
    - All tubes in the left trigger logic was timed against a single big
    tube in the right array, using discriminators masks, and adjusting the
    programmable delay. Still need to do the small tubes in the right
    array. This is a prequisite for the final trigger adjustment.
    
    - As described on one of the mailing lists Peter has commited
    a new and slim version of the Tof Online Monitor. A single detector 
    now has only a few 
     A few suggestion
    -- Put the pedestal distributions on the same page as the
    masetr summaries.
    -- The dots on the 2d hists are way too big. Use box and/or 
    logaritmic representation
    
    As we are getting these slimmed down monitor plots it is also
    necs. to have a brief description e.g. in the form of a 'default plots'
    for normal conditions with annotations.
    
    having used it a bit.
    i) 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Jul 20 2000 - 00:30:44 EDT